OpenStudy (anonymous):

Thor invests $3000 in an account, with interest compounded continuously. If his investment doubles in value after 7.2 years, what interest rate is he earning? [Use the formla A = Pe rt and round answer to the nearest tenth.]

OpenStudy (anonymous):

e^r*7.2=2 7.2r=In(2) r=In(2)/7.2